OpenStudy (anonymous):

48xy (y is too the 4th power) - 72xy (y is too the 2nd power) ------------ (divided by) 12xy(y is too the 2nd power)

OpenStudy (anonymous):

\[48xy ^{4}-72xy ^{2} \div 12xy ^{2}\] Here is a clearer version of the equation! (:

OpenStudy (anonymous):

\[\frac{ 48xy^4-72xy^2 }{ 12xy^2 }\]did you notice anything about the top? you can pull out a y^2 from both of those numbers so lets do it\[\frac{ y^2(48xy^2-72x) }{ 12xy^2 }=\frac{ 48xy^2-72x }{ 12x }\]since like terms on top and bottom cancel out, and now look again, we have more common factors! we can pull an x from the top, as well as a 12 so let's do that\[\frac{ 12x(4y^2-6) }{ 12x }=4y^2-6\]and that's a much simpler, and equivalent way to right that problem

OpenStudy (anonymous):

Thank you! This help means a lot to me because math really isn't my strong subject haha! (: